OpenStudy (anonymous):

how to find the indefinite integral of (x^2-1)/x

OpenStudy (anonymous):

Use a calculator noob... (x^2-1)/x can be rewritten as (x^2)/x -1/x split up the integral \[\int\limits_{?}^{?}x - \int\limits_{?}^{?}1/x = x^2/2 + \ln \left| x \right| + c\]